(1 point) Consider the initial value problem my+cy+ky F(t), y(0) = 0, y(0) = 0 modeling the motion of a spring-mass-dashpo
Answer #1

where Given initial value problem is my Acy + ky = f(1)_700)=b, yro)=0 m = 2 kg c=8kg, K: 80N/m und Elt) = 60 cosst N - Gi4 3-6 15 1r+6) 18-6)(8+6) cosst - 15 lote) g2-36 oset = 15 2 18+6) cosst -64-3€ = cosst 15 2 (8+6) -100 200 = -15 18 losst+ 6How differentiate og? 0 word to y rozēt [ geesot +95in£t 1+ ē7-gesinett Ga6 teset) - 3 / - C4 cos 87 - 48110 3+1 using ylo)=